# Automotive Speed Estimation: Sensor Types and Error Characteristics from OBD-II to ADAS

### v1

@@ I Introduction
Many complementary systems, such as navigation and driver assistance solutions added post-manufacture, access essential vehicular data through the On-Board Diagnostic II (OBD-II) port [3]. As illustrated in Figure 1, the OBD-II interface connects to the vehicle’s Controller Area Network (CAN) bus, enabling communication with subsystems like the Electronic Control Units (ECUs), Anti-lock Braking System (ABS), Instrument Cluster, and Adaptive Cruise Control (ACC) modules, to name a few. This standardized interface provides critical parameters, including engine RPM and wheel speed, facilitating seamless integration with internal and external systems. However, the sensors and techniques used to derive speed measurements can differ significantly between land vehicles. These differences can introduce variations in error characteristics, which can profoundly impact the reliability and accuracy of downstream systems relying on this data.
@@
Wheel speed, in particular, plays a critical role in enhancing Inertial Navigation Systems (INS) and enabling the Reduced Inertial Sensor System (RISS), a simplified INS variant that depends entirely on external speed inputs for operation [4]. The quality of wheel speed data is essential in GNSS-denied environments or, in the case of modern systems, signals of opportunity (SOPs)-denied conditions, where errors in proprioceptive measurements can propagate and degrade dead-reckoning performance. Furthermore, a limited understanding of sensor types, error characteristics, and data formats can compromise the accuracy and reliability of complementary systems.
@@
    Comprehensive overview of automotive speed sensing techniques: We provide a detailed analysis of automotive speed sensing methodologies, offering insights into how vehicle speed is estimated using classical and modern sensor technologies.
@@
The remainder of this paper is organized as follows: Section II reviews related work on ground vehicle speed sensing and its applications in navigation. Then, Section III explores OBD-II communication, data formatting, and speed derivation for various vehicle configurations, along with an overview of WSS. Following this, Section IV presents experimental results, including sensor identification, error variance analysis, and the integration of stereo cameras for complementary speed estimation. Finally, Section V concludes the paper and outlines future research directions.
@@ II Related Work
In addition to the often omitted derivation process, the navigation and instrumentation literature offers limited clarity on the specific sensors automotive suppliers use for speed estimation. Researchers often overlook the type of speed sensors employed in their test vehicles, commonly employ the general term odometer which covers a wide variety of instruments, and frequently rely on traditional Inertial Measurement Unit (IMU) and GNSS error modeling techniques for sensor fusion [4, 12, 13, 14]. While these approaches have been widely used, they may not guarantee optimal estimation of vehicle speed. Therefore, identifying and characterizing vehicular speed sensors is critical for designing robust and effective navigation systems. Figure 2 categorizes vehicle speed-sensing technologies into two primary groups: built-in diagnostic data and alternative sensors and systems. The first group includes OBD-II-derived speed measurements and CAN bus-derived data from Vehicle Speed Sensors (VSS) and Wheel Speed Sensors (WSS), both of which are widely integrated into vehicles’ onboard systems. The second group consists of external sensors, such as non-contact optical sensors, DMIs, and perception-based systems like cameras, LiDARs, and RADARs. Cameras and RADARs, in particular, are increasingly utilized not only for ADAS but also for forward velocity estimation. Recent studies have demonstrated the effectiveness of cameras for ego-motion tracking [15] and RADARs for ego-velocity estimation [16]. These perception-based approaches offer practical and cost-effective alternatives, particularly in scenarios where traditional automotive speed sensors are limited or unavailable.
@@ III OBD-II Vehicle Speed Derivation
The command 010D represents a request for the current vehicle speed (PID 0x0D). The response 41 0D 3C indicates the vehicle speed as 0x3C, which is 60 $km/h$ when converted to decimal. In other words, the reported speed through OBD-II interfacing is:
$V_{\text{v}}^{\text{OBD}}=\left\lfloor\hat{V}\right\rceil$ (1)
@@ III-A Transmission-based Vehicle Speed Determination
The estimation of vehicle speed ($V_{\text{v}}$) using transmission data is based on the relationship between the engine speed ($N_{e}$), expressed in revolutions per minute (RPM), and key drivetrain parameters. A critical component of this relationship is the gear ratio, which defines how the rotational speed of the engine is modified by the transmission and differential before reaching the wheels.
The gear ratio determines the relationship between the rotational speeds of the input (engine) and output (wheels). For example, as illustrated in Figure 3, a driving gear with 8 teeth and a driven gear with 16 teeth results in a gear ratio of 2:1, meaning the output rotates at half the speed of the input. This adjustment is essential for controlling the vehicle’s speed and torque dynamics.
@@
In non-ABS vehicles, where speed is derived from the engine’s VSS (Figure 4), the calculation for speed in $km/h$ is given by the following formula:
$V_{\text{v}}=\left\lfloor\left({\frac{N_{e}\cdot\pi\cdot r_{w}\cdot 3.6}{30\cdot i_{x}\cdot i_{0}}}\right)\right\rceil$ (2)
where $N_{e}$ is the engine speed, while $i_{x}$ and $i_{0}$ represent the dimensionless gear ratio of the engaged gear and the differential gear ratio, respectively. The wheel radius ($r_{w}$), expressed in meters ($m$), converts the rotational speed of the wheels into linear speed. This calculation assumes a slip-free mechanical connection between the engine and wheels, ensuring precise power transfer and accurate speed estimation.
@@ III-B ABS-based Vehicle Speed Determination
The ABS WSS produces output signals as the wheel turns, i.e. wheel ticks. These signals are then transmitted to the vehicle’s ECU or a specific ABS control module. The main role of the ECU or ABS module is to interpret these signals to determine each wheel’s speed.
@@
$V_{\text{w}}=\frac{2\pi\cdot r_{w}\cdot N_{p}}{N_{t}\cdot\Delta t}$ (3)
@@
$V_{\text{v}}=\left\lfloor\left(\frac{V_{\text{w}_{1}}+V_{\text{w}_{2}}+V_{\text{w}_{3}}+V_{\text{w}_{4}}}{4}\right)\right\rceil$ (4)
@@ III-C Wheel Speed Sensor Types
- Limited diagnostic - More susceptible to EMI
capabilities
@@ IV-A Experimental Setup
The experiments were conducted using an ABS-equipped Toyota Sienna minivan, with the logging setup briefly depicted in Figure 6. The dataset collected during these experiments has been made publicly available. [^1] The dataset can be accessed at: https://github.com/hanymragab/navinst19
@@ IV-B1 Identification and Error Modeling of Automotive Speed Sensors
Figure 7 shows the results for the data collected in the cities of Kingston and Toronto that were used to fit the error variance. Upon initial data inspection, the quantity of interest that was found to impact measurement accuracy was the speed of travel. This is a clear indication that the sensor used is $passive$ in its nature. Moreover, the reason why the 0-4 $km/h$ speed interval was omitted in the variance analysis is simply due to the fact that PWSS may not report speeds lower than around 3 $km/h$ as shown in Figure 8.
@@
Driving patterns vary significantly depending on the environment, with vehicles typically traveling slower in dense urban areas and downtown cores compared to highways and other parts of the road network. Figure 9 illustrates this trend, showing that the probability of a car traveling at speeds between 0–4 $km/h$ is 35–45% for trajectories T19.1 and K19.1, which traverse urban and downtown cores. In contrast, for T19.2, which includes a significant portion of highways, this probability drops below 30%. In dense, GNSS-denied urban areas, evaluating perception systems for speed estimation becomes particularly relevant.
@@ IV-B2 OBD-II-derived speed vs Stereo VO-derived speed
To evaluate the performance of speed estimates across different intervals, we compare OBD-II-derived vehicle speeds with results from a stereo visual odometry (Stereo VO) routine and an enhanced Stereo VO algorithm incorporating semantic segmentation-based outlier rejection (Stereo VO-SS-OR), as introduced in our earlier work [18]. Figure 10 reveals significant variations in the accuracy of speed measurements. In the lower speed range (0–4 $km/h$), Stereo VO-SS-OR outperforms the PWSS, achieving a lower root mean square error (RMSE), whereas PWSS exhibits higher RMSE in this range, unlike at higher speed intervals.
@@
The results demonstrate variability in the performance of Stereo VO under different scenarios. Figure 11 showcases the potential of Stereo VO, particularly the enhanced version with SS-OR, to improve vehicular speed estimation when PWSS fails to report speed. The comparison includes speed estimates from Stereo VO, both with and without the SS-OR enhancement, alongside OBD-II-derived speeds and the NovAtel reference. The results demonstrate that integrating SS-OR into Stereo VO significantly improves ego-motion estimation accuracy, leading to more precise speed estimation.
@@
On the other hand, Figure 12 illustrates the challenges faced in low-traffic environments dominated by static objects, which significantly affect the approach’s effectiveness. In such scenarios, Stereo VO generates noisier and less accurate speed estimates compared to OBD-II-derived speeds. The performance of Stereo VO-SS-OR deteriorates further at higher travel speeds, primarily due to baseline discrepancies between frames combined with framerate limitations. Additionally, the abundance of parked cars (as highlighted in the figure), often being misclassified as dynamic within the SS-OR routine,
@@
disrupts temporal feature tracking and impairs accurate ego-motion estimation. Given these differences, implementing a fusion scheme is essential to selectively weight measurements based on WSS speed data, enabling a more accurate estimation of forward speed. Additionally, this highlights the critical need for integrating object tracking or supplementary sensors, such as RADAR sensors to improve stationary and dynamic object classification, enhancing the overall performance of the SS-OR framework.
@@ V Conclusion
In conclusion, this paper sheds light on a critical yet underexplored aspect of navigation and instrumentation: the diverse sensors and methods of deriving automotive speed measurements from the OEM perspective. By identifying and analyzing these approaches, we demonstrated the strengths and limitations of using OBD-II-derived speed data for land vehicle navigation. Through experiments conducted on three extended trajectories, we identified the wheel speed sensor type employed and showcased the advantages of using alternative ADAS sensors, such as stereo cameras or radars, for speed estimation.

### v2

@@ I Introduction
Many complementary systems, such as navigation and driver assistance solutions added post-manufacture, access essential vehicular data through the On-Board Diagnostic II (OBD-II) port [3]. As illustrated in Fig. 1, the OBD-II interface connects to the vehicle’s primary network, typically the CAN bus, or other protocols depending on the make and model. It allows communication with subsystems such as the Electronic Control Units (ECUs), Anti-lock Braking System (ABS), Instrument Cluster, and Adaptive Cruise Control (ACC) modules, to name a few. This standardized interface provides critical parameters, including engine RPM and wheel speed, facilitating seamless integration with internal and external systems. However, the sensors and techniques used to derive speed measurements can differ significantly between land vehicles. These differences can introduce variations in error characteristics, which can profoundly impact the reliability and accuracy of downstream systems relying on this data.
@@
Wheel speed, in particular, plays a critical role in enhancing Inertial Navigation Systems (INS) and enabling the Reduced Inertial Sensor System (RISS), a simplified INS variant that predominantly depends on external speed inputs for operation [4]. The quality of wheel speed data is essential in GNSS-denied environments or, in the case of modern systems, signals of opportunity (SOPs)-denied conditions, where errors in proprioceptive measurements can propagate and degrade dead-reckoning performance. Furthermore, a limited understanding of sensor types, error characteristics, and data formats can compromise the accuracy and reliability of complementary systems. The key contributions of this paper are:
@@
    Comprehensive overview of automotive speed sensing: We analyze methods for estimating vehicle speed, highlighting classical and modern sensor technologies.
@@
The remainder of this paper is organized as follows: Section II reviews related work on ground vehicle speed sensing and its applications in navigation. Then, Section III explores OBD-II communication, data formatting, and speed derivation for various vehicle configurations, along with an overview of wheel speed sensors (WSS). Following this, Section IV presents experimental results, including sensor identification, error variance analysis, and the integration of stereo cameras for complementary speed estimation. Finally, Section V concludes the paper and outlines future research directions.
@@ II Related Work
In addition to the often omitted derivation process, the navigation and instrumentation literature offers limited clarity on the specific sensors automotive suppliers use for speed estimation. Researchers often overlook the type of speed sensors employed in their test vehicles, commonly employ the general term odometer which covers a wide variety of instruments, and frequently rely on traditional Inertial Measurement Unit (IMU) and GNSS error modeling techniques for sensor fusion [4, 12, 13, 14]. While these approaches have been widely used, they may not guarantee optimal estimation of vehicle speed. Therefore, identifying and characterizing vehicular speed sensors is critical for designing robust and effective navigation systems. Fig. 2 categorizes vehicle speed-sensing technologies into two primary groups: built-in diagnostic data and alternative sensors and systems. The first group includes OBD-II-derived speed measurements and CAN bus-derived data from Vehicle Speed Sensors (VSS) and WSS, both of which are widely integrated into vehicles’ onboard systems. The second group consists of external sensors, such as non-contact optical sensors, DMIs, and perception-based systems like cameras, LiDARs, and RADARs. Cameras and RADARs, in particular, are increasingly utilized not only for ADAS but also for forward velocity estimation. Recent studies have demonstrated the effectiveness of cameras for ego-motion tracking [15] and RADARs for ego-velocity estimation [16]. These perception-based approaches offer practical and cost-effective alternatives, particularly in scenarios where traditional automotive speed sensors are limited or unavailable.
@@ III OBD-II Vehicle Speed Derivation
The command 010D represents a request for the current vehicle speed (PID 0x0D). The response 41 0D 3C indicates the vehicle speed as 0x3C, which is 60 $km/h$ when converted to decimal [^1] Automakers may configure ECUs to report diagnostic data in region-specific units. For example, vehicle speed may be in miles per hour ($mph$) for the USA or in kilometers per hour ($km/h$) for Canada and Europe.. In other words, the reported speed through OBD-II interfacing is:
$V_{\text{v}}^{\text{OBD}}=\left\lfloor\hat{V}\right\rceil.$ (1)
@@ III-A Transmission-based Vehicle Speed Determination
The estimation of vehicle speed ($V_{\text{v}}$) from transmission data relies on the relationship between engine speed ($N_{e}$) in RPM and drivetrain parameters, especially gear ratio. The gear ratio governs the conversion of engine speed (input) to wheel speed (output). For example, in Fig. 3, a driving gear with 8 teeth and a driven gear with 16 teeth creates a 2:1 ratio, where the output rotates at half the input speed. This adjustment is crucial for controlling speed and torque. The gearbox alters engine speed based on the engaged gear ratio ($i_{x}$), which changes with each gear shift, while the differential uses its gear ratio ($i_{0}$) to balance wheel speeds, critical for smooth turning.
@@
In non-ABS vehicles, where speed is derived from the engine’s VSS (see Fig. 4), the calculation for speed in $m/s$ is given by the following formula:
$V_{\text{v}}=\left\lfloor\left({\frac{N_{e}\cdot\pi\cdot r_{w}}{30\cdot i_{x}\cdot i_{0}}}\right)\right\rceil,$ (2)
$N_{e}$ represents the engine speed, typically measured by a crankshaft position (CKP) sensor or an engine speed sensor. The wheel radius ($r_{w}$), measured in meters ($m$), converts the rotational speed of the wheels into linear speed. This calculation assumes a slip-free mechanical connection between the engine and wheels, ensuring efficient power transfer and accurate speed estimation.
@@ III-B ABS-based Vehicle Speed Determination
The ABS WSS, depicted in Fig. 5, produces output signals as the wheel turns, i.e. wheel ticks. These signals are then transmitted to the vehicle’s ECU or a specific ABS control module. The main role of the ECU or ABS module is to interpret these signals to determine each wheel’s speed.
@@
$V_{\text{w}}=\frac{2\pi\cdot r_{w}\cdot N_{p}}{N_{t}\cdot\Delta t},$ (3)
@@
$V_{\text{v}}=\left\lfloor\left(\frac{V_{\text{w}_{1}}+V_{\text{w}_{2}}+V_{\text{w}_{3}}+V_{\text{w}_{4}}}{4}\right)\right\rceil,$ (4)
@@ III-C Wheel Speed Sensor Types
- Limited diagnostic - More susceptible to
capabilities electromagnetic interference
@@ IV-A Experimental Setup
The experiments were conducted using an ABS-equipped Toyota Sienna minivan, with the logging setup briefly depicted in Fig. 6. The dataset collected during these experiments has been made publicly available [^2] The dataset can be accessed at: https://github.com/hanymragab/navinst19.
@@ IV-B1 Identification and Error Modeling of Automotive Speed Sensors
Fig. 7 shows the results for the data collected in the cities of Kingston and Toronto that were used to fit the error variance. Upon initial data inspection, the quantity of interest that was found to impact measurement accuracy was the speed of travel. This is a clear indication that the sensor used is $passive$ in its nature. Moreover, the reason why the 0-4 $km/h$ speed interval was omitted in the variance analysis is simply due to the fact that PWSS may not report speeds lower than around 3 $km/h$ as shown in Fig. 8.
@@
Driving patterns vary significantly depending on the environment, with vehicles typically traveling slower in dense urban areas and downtown cores compared to highways and other parts of the road network. Fig. 9 illustrates this trend, showing that the probability of a car traveling at speeds between 0–4 $km/h$ is 35–45% for trajectories T19.1 and K19.1, which traverse urban and downtown cores. In contrast, for T19.2, which includes a significant portion of highways, this probability drops below 30%. In dense, GNSS-denied urban areas, evaluating perception systems for speed estimation becomes particularly relevant.
@@ IV-B2 OBD-II-derived speed vs Stereo VO-derived speed
To evaluate the performance of speed estimates across different intervals, we compare OBD-II-derived vehicle speeds with results from a stereo visual odometry (Stereo VO) routine and an enhanced Stereo VO algorithm incorporating semantic segmentation-based outlier rejection (Stereo VO-SS-OR), as introduced in our earlier work [18]. Fig. 10 reveals significant variations in the accuracy of speed measurements. In the lower speed range (0–4 $km/h$), both Stereo VO algorithms outperform the OBD-derived PWSS speed measurements, with Stereo VO-SS-OR achieving the lowest root mean square error (RMSE), unlike at higher speed intervals.
@@
Fig. 11 presents a comparison of speed estimates from Stereo VO, with and without the SS-OR enhancement, alongside OBD-II-derived speeds and the NovAtel reference. It highlights the variability in Stereo VO performance across different scenarios and showcases the capability of the SS-OR-enhanced version to improve vehicular speed estimation, particularly when PWSS fails to provide data. The integration of SS-OR significantly enhances ego-motion estimation accuracy, resulting in more consistent and precise speed measurements, especially at lower speeds of travel in highly dynamic environments.
@@
On the other hand, Fig. 12 illustrates the challenges faced in low-traffic environments dominated by static objects, which significantly affect the approach’s effectiveness. In such scenarios, Stereo VO generates noisier and less accurate speed estimates compared to OBD-II-derived speeds. The performance of Stereo VO-SS-OR deteriorates further at higher travel speeds, primarily due to baseline discrepancies
@@
between frames combined with framerate limitations. Additionally, the abundance of parked cars (as highlighted in the figure), often being misclassified as dynamic within the SS-OR routine, disrupts temporal feature tracking, and impairs accurate ego-motion estimation. Given these differences, implementing a fusion scheme is essential to selectively weight measurements based on WSS speed data, enabling a more accurate estimation of forward speed. Additionally, this highlights the critical need for integrating object tracking or supplementary sensors, such as RADAR sensors to improve stationary and dynamic object classification, enhancing the overall performance of the SS-OR framework.
@@ V Conclusion
In conclusion, this paper sheds light on a critical yet underexplored aspect of navigation and instrumentation: the diverse sensors and methods of deriving automotive speed measurements from the OEM perspective. By identifying and analyzing these approaches, we demonstrated the strengths and limitations of using OBD-II-derived speed data for land vehicle navigation. Through experiments conducted on three extended trajectories, we identified the wheel speed sensor type employed and showcased the advantages of using alternative ADAS sensors, such as front-facing stereo cameras, for speed estimation.